# Quantum property testing in sparse directed graphs

### v2

@@
Quantum property testing in sparse directed graphs
Simon Apers Email: {apers, magniez, szabo}@irif.fr. Research supported in part by the ERC Advanced Grant PARQ (grant agreement No 885394), the European QuantERA project QOPT (ERA-NET Cofund 2022-25), the French PEPR integrated projects EPiQ (ANR-22-PETQ-0007) and HQI (ANR-22-PNCQ-0002), and the French ANR project QUOPS (ANR-22-CE47-0003-01). Université Paris Cité, CNRS, IRIF, Paris, France Frédéric Magniez11footnotemark: 1 Université Paris Cité, CNRS, IRIF, Paris, France Sayantan Sen Email: sayantan789@gmail.com. Research supported by the National Research Foundation, Singapore through the National Quantum Office, hosted in A*STAR, under its Centre for Quantum Technologies Funding Initiative (S24Q2d0009). Centre for Quantum Technologies, National University of Singapore, Singapore Dániel Szabó11footnotemark: 1 Université Paris Cité, CNRS, IRIF, Paris, France
@@ 1.1 Context
Graphs are of paramount importance for instance when it comes to understanding large datasets, since they provide a natural way to represent and analyze complex relationships inside datasets. Goldreich, Goldwasser, and Ron [goldreich1998property] were the first to consider graphs in the context of property testing. Formally, given some form of query access to an unknown graph $G$ on $N$ vertices, and a property $\mathcal{P}$ of interest, the goal is to distinguish with high probability if $G$ satisfies the property $\mathcal{P}$, or whether it is “far” from all graphs that satisfy $\mathcal{P}$, with a suitable notion of farness. In [goldreich1998property], the “dense” model was considered, where a graph is accessed through adjacency queries: for a pair of vertices $(u,v)$, the query reveals whether $(u,v)$ is an edge in the graph. In this model, a graph $G$ is $\varepsilon$-far from satisfying $\mathcal{P}$ if one needs to add or remove at least $\varepsilon N^{2}$ edges of $G$ to obtain a graph that satisfies $\mathcal{P}$.
In a later work, Goldreich and Ron [goldreich2002property] introduced the “bounded-degree” model for testing sparse graphs, focusing on the properties of bipartiteness and expansion. In this model, a $d$-bounded degree graph $G$ with $N$ vertices is accessed by performing neighbor queries: for a vertex $v$ and an integer $i\in[d]$, the query $(v,i)$ returns either the $i$-th neighbor of $v$, or some special symbol if $v$ has less than $i$ neighbors. The graph $G$ is said to be $\varepsilon$-far from some property $\mathcal{P}$, if one needs to add or delete at least $\varepsilon dN$ edges of $G$ to obtain a graph that satisfies $\mathcal{P}$. Over the last two decades, there has been a significant number of works in this model, and we refer the interested reader to the books by Goldreich [goldreich2017introduction] and Bhattacharyya and Yoshida [bhattacharyya2022property] and several surveys [DBLP:journals/eatcs/Fischer01, DBLP:journals/fttcs/Ron09, Czumaj2010, DBLP:journals/siamdm/RubinfeldS11].
@@ 1.3 Related works on collision finding
The problem of collision finding is a ubiquitous problem in the field of algorithm theory with wide applications in cryptography. Here, given a sequence $s$ of $N$ integers, the goal is to find a duplicate in $s$. If one has the guarantee that $\Theta(N)$ elements of the sequence are duplicated, which is the case for instance when the sequence consists of uniformly random integers from $[N]$, it is well-known that classically $\Theta(\sqrt{N})$ queries are necessary and sufficient due to the birthday paradox. In the quantum model, this can be solved with query complexity $\Theta(N^{1/3})$ by the algorithm of Brassard, Høyer and Tapp [BHT]. The matching lower bound was first stated for a specific set of hard instances known as $2$-to-$1$ (i.e., each integer appears exactly twice or not at all) by Aaronson and Shi [AaronsonShi]. For some constant integer $k\geq 3$, those results can be further extended to finding $k$-collisions in a random input with suitable alphabet size, so that it contains $\Theta(N)$ $k$-duplicates with high probability. The classical query complexity for this problem is $\Theta(N^{1-1/k})$ [hellweg2013property, peng2023optimal], and quantumly it is $\Theta\left(N^{\frac{1}{2}\left(1-\frac{1}{2^{k}-1}\right)}\right)$ [zhandry_collision]. The situation is more complex for non-random inputs.
Remarkably, the complexity of testing $k$-collision-freeness (i.e., the absence of $k$-collisions) is harder to settle on the lower bound side than the finding version. In this work, we are going to focus on the hardness of distinguishing inputs that have linearly many collisions from those that do not have any. For $k=2$, the two problems have the same complexity, since intuitively the only way to distinguish is to find a collision. This can be formalized easily in the classical case. Quantumly, this is more challenging, but the lower bound in [AaronsonShi] proved the hardness of distinguishing between $2$-to-$1$ instances and ones with no duplicate.
@@ 1.4 Our results
Theorem 1.1 (Restated in Theorem 3.3).
@@
Theorem 1.2 (Corollary of Theorem 1.3).
@@
For proving graph property testing lower bounds, both the classical works of [hellweg2013property] and [peng2023optimal] prove collision testing lower bounds using the proportional moments technique of [raskhodnikova2009strong]. At the heart of this technique is a construction of two positive integer random variables, $X_{1}$ and $X_{2}$, with different expectations but with the following conditions on the first $k-1$ moments: $\mathop{\mathbb{E}}[X_{1}]/\mathop{\mathbb{E}}[X_{2}]=\mathop{\mathbb{E}}[X^{2}_{1}]/\mathop{\mathbb{E}}[X^{2}_{2}]=\ldots=\mathop{\mathbb{E}}[X^{k-1}_{1}]/\mathop{\mathbb{E}}[X^{k-1}_{2}]$. Such a construction leads to a randomized query complexity lower bound of $\Omega(N^{1-\frac{1}{k}})$ for various property testing problems such as $k$-collision-freeness [peng2023optimal]. Having a quantum version of this technique has been identified as an important open problem [ambainis2016efficient], since this could be used to pave the way to stronger quantum lower bounds in related settings. We modestly made progress to this quest for the special case of testing $k$-collision-freeness.
@@
Theorem 1.3 (Restated in Theorem 4.1).
@@
Theorem 1.4 (Restated in Theorem 5.1).
@@ 1.5.1 Subgraph-finding algorithm
We start by describing how to prove the upper bound result of Theorem 1.1 for testing $k$-source-subgraph-freeness. We view the problem as a generalization of the problem of finding $k$-collisions and adapt an existing quantum algorithm for the latter problem. In [zhandry_collision], an algorithm is given for finding $k$-collisions in length-$N$ sequences of integers that contain $\Omega(N)$ $k$-collisions (e.g. $k$-to-$1$, or random sequence with appropriate parameters). Their algorithm generalizes the well-known collision finding algorithm of [BHT]. On a high level, the [zhandry_collision] algorithm first finds several $2$-collisions using Grover search like in [BHT], extends some of them to $3$-collisions in a similar way, and so on until a $k$-collision is found.
@@
When we make the second step of turning to testing of subgraph-freeness, we need to prove a variant of this statement: if an $N$-vertex graph $G$ is far from $H$-freeness (for some constant-sized subgraph $H$) then it contains $\Omega(N)$ many “source-disjoint” $H$-subgraphs. This means that there are $\Omega(N)$ many such $H$-subgraphs in $G$ that the set of vertices in the source components of each $H$-subgraph are disjoint. We prove this fact in Proposition 3.2 and this allows us to further generalize the approach of [zhandry_collision]: first find several partial solutions where only a few source components of an $H$-subgraph are explored, and gradually extend these (using Grover search coupled with constant-depth BFS) until a complete $H$-subgraph is found.
@@ 1.5.2 Collision-freeness lower bound
Now we will discuss our approach to proving the lower bounds of collision-freeness (Theorem 1.2) and $k$-source-subgraph-freeness (Theorem 1.3). We first give a simple reduction from $k$-collision-freeness to $k$-star-freeness, which is a special case of $k$-source-subgraph-freeness. This way, it is enough to prove a lower bound on testing $k$-collision-freeness, and it implies the same result on testing $k$-source-subgraph-freeness. Since our lower bound approach crucially depends on the (dual) polynomial method, let us start by briefly discussing it.
@@
The polynomial method can be written in the form of a linear program, of which one can take the dual. By weak LP-duality, when using this dual characterization for proving a lower bound on function $f$, one needs to provide a “witness” of the approximating polynomial’s high degree, say $\Delta$. This witness is called the dual polynomial $\psi$ and, in the easiest case of total Boolean [^1] We use $\{-1,1\}$ where $-1$ corresponds to the “true value”. functions $f:\{-1,1\}^{n}\to\{-1,1\}$, it needs to have three properties:
@@
    Pure high degree $\Delta$: $\sum_{x}p(x)\psi(x)=0$, for every polynomial $p$ with degree $<\Delta$,
@@
$F(s)=\begin{cases}-1&\text{if no integer occurs at least }k\text{ times in }s,\\ 1&\text{if more than }\gamma R\text{ distinct integers occur at least }k\text{times in }s,\\ \text{undefined}&\text{otherwise}.\end{cases}$
@@
Moreover, the symmetries of $f$ allow the extension of the initial function $f$ from the very restricted set of binary inputs correspond to valid strings, to the more general set of binary inputs of Hamming weight $N$ [ambainis2016efficient]. With further technicalities one can also extend $f$ to all binary inputs of Hamming weight at most $N$ [BT20]. This is fundamental because instead of being forced to zero out the dual polynomial outside the domain of $f$, we only need to do so on inputs of Hamming weight higher than $N$. Using the symmetry of $f$, it can be shown that a lower bound on this modified, Boolean version implies a lower bound on the original $k$-collision problem.
@@
    Pure high degree $\Delta$: $\sum_{x}p(x)\psi(x)=0$, for every polynomial $p$ with degree $<\Delta$;
@@
Coming back now to the definition of our Boolean function $f$, one can rewrite it as a composition of simpler functions: $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$, where by composition we mean $(g\circ h)(x)=g(h(x_{1}),\dots,h(x_{n}))$ (where $x=(x_{1},\dots,x_{n})$ and each $x_{i}$ is a binary vector of appropriate dimension) and the domain is restricted to bit strings of Hamming weight at most $N$. Note that here, $x_{j}=(x_{1,j},x_{2,j},\ldots,x_{N,j})$. Above, $\mathrm{THR}_{N}^{k}$ is the threshold function: it is $-1$ if the input bitstring contains at least $k$ many $-1$ (true) values, and is $1$ otherwise; and $\mathrm{GapOR}_{R}^{\gamma}$ is the gap version of $\mathrm{OR}$, which is $1$ if the input is all-1, $-1$ if the input contains at least $\gamma R$ many $-1$ values, and is undefined otherwise.
In order to give a dual polynomial for this composed function, we start from a dual polynomial for each part of the composition ($\phi$ and $\psi$), which were already given in [polynomial_strikes_back] (in different contexts). Then we use a known way [ShiZ09, Lee09, Sher13] of composing dual polynomials called the dual block composition, which provides a nearly good dual polynomial $\phi\star\psi$ for $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$. Indeed, by construction, the normalization (ii) and the pure high degree (iii) are guaranteed. However, the issue is that it is not 00 on bitstrings of large Hamming weight thus (iv) is not satisfied, and and the high correlation (i’) still has to be proved.
To fix (iv), we use another result of [polynomial_strikes_back] which provides another dual polynomial $\zeta$, that is close to $\phi\star\psi$ and that is 00 on inputs having Hamming weight larger than $N$. Also, it only changes the pure high degree by a polylogarithmic factor. Now the only remaining task is to prove a large enough correlation (i’) of $\phi\star\psi$ and $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$ so that $\zeta$ still has high enough correlation. This high correlation proof (Lemma 4.18) is the most technical part of our paper.
High correlation: proof of Lemma 4.18
@@
In the proof of this lemma we use Proposition 4.17 that is a more general statement of some techniques used in several proofs of [polynomial_strikes_back]. But then we need to diverge from their proof because it crucially relies on a certain one-sided error property (in the sense of [polynomial_strikes_back, Lemma 6.11]) of the inner function of the composition which is OR function in their case. Our inner function is the threshold function which does not satisfy this property, so we have to use some other properties of the dual polynomials in our proof. This different proof technique could be a step towards obtaining a more general lower bound technique.
@@
Let us now discuss our approach to proving the linear lower bound on quantum query complexity of testing $3$-colorability (Theorem 1.4). Before proceeding to present our approach, let us briefly discuss the classical lower bound of testing $3$-colorability. To prove the classical lower bound of $3$-colorability, the authors in [bogdanov2002lower] first studied another problem called E$(3,c)$LIN-2, a problem related to deciding the satisfiability of a system of linear equations. More formally, E$(3,c)$LIN-2 considers a system of linear equations modulo $2$, where each equation has $3$ variables and every variable appears in at most $c$ equations. Given such a system of linear equations, the goal is to distinguish if it is satisfiable, or at least some suitable fraction of the equations need to be modified to satisfy it. [bogdanov2002lower] proved that $\Omega(N)$ classical queries to the system of linear equations are necessary for testing E$(3,c)$LIN-2.
After this, they designed a reduction from E$(3,c)$LIN-2 to $3$-colorability such that satisfying instances of E$(3,c)$LIN-2 are reduced to $3$-colorable graphs, and far from satisfiable instances of E$(3,c)$LIN-2 are mapped to far from $3$-colorable graphs. Combining these two arguments, the authors in [bogdanov2002lower] proved that $\Omega(N)$ classical queries are necessary for testing $3$-colorability for bounded degree graphs.
The authors in [bogdanov2002lower] used Yao’s minimax method to prove the linear lower bound in testing E$(3,c)$LIN-2. In particular, they designed two distributions $D_{\mathrm{yes}}$ and $D_{\mathrm{no}}$ such that the systems of linear equations in $D_{\mathrm{yes}}$ are satisfiable, whereas the systems of linear equations in $D_{\mathrm{no}}$ are far from being satisfiable. A crucial ingredient of their lower bound proof is a construction of a system of linear equations (represented as a matrix) that are far from being satisfiable, but any $\delta N$ rows of the matrix are linearly independent. Hence, any subset of $\delta N$ entries of the matrix will look uniformly random, and therefore hard to distinguish from a satisfiable instance.
It is a known fact that distinguishing between a uniformly random string and a $\ell$-wise independent string, for an appropriate integer $\ell$, is hard for quantum algorithms (see e.g. [apers2022quantum]). Using this result, we can construct suitable hard instances for E$(3,c)$LIN-2 such that testing E$(3,c)$LIN-2 remains maximally hard (requires $\Omega(N)$ queries) for any quantum algorithm. Combining this hardness result with the reduction from E$(3,c)$LIN-2 to $3$-colorability, we finally prove that $\Omega(N)$ quantum queries are necessary for testing $3$-colorability. We formally prove this in Section 5.

### v3

@@
Quantum property testing in sparse directed graphs  [^1] A preliminary version of this work appeared in Approximation, Randomization, and Combinatorial Optimization. Algorithms and Techniques (APPROX/RANDOM), volume 353, pages 32:1–32:24. Schloss Dagstuhl – Leibniz-Zentrum für Informatik, 2025 [AMSS25_random].
Simon Apers Email: {apers, magniez, szabo}@irif.fr. Research supported in part by the ERC Advanced Grant PARQ (grant agreement No 885394), the European QuantERA project QOPT (ERA-NET Cofund 2022-25), the French PEPR integrated projects EPiQ (ANR-22-PETQ-0007) and HQI (ANR-22-PNCQ-0002), and the French ANR project QUOPS (ANR-22-CE47-0003-01). Université Paris Cité, CNRS, IRIF, Paris, France Frédéric Magniez22footnotemark: 2 Université Paris Cité, CNRS, IRIF, Paris, France Sayantan Sen Email: sayantan789@gmail.com. Research supported by the NRF Investigatorship award (NRF-NRFI10-2024-0006) and CQT Young Researcher Career Development Grant (25-YRCDG-$S$). Centre for Quantum Technologies, National University of Singapore, Singapore Dániel Szabó22footnotemark: 2 Research supported in part by the German Federal Ministry of Research, Technology and Space (QuSol, 13N17173). Université Paris Cité, CNRS, IRIF, Paris, France Ruhr-Universität Bochum, Bochum, Germany
@@ 1.1 Context
Graphs are of paramount importance for instance when it comes to understanding large datasets, since they provide a natural way to represent and analyze complex relationships inside datasets. Goldreich, Goldwasser, and Ron [GGR98] were the first to consider graphs in the context of property testing. Formally, given some form of query access to an unknown graph $G$ on $N$ vertices, and a property $\mathcal{P}$ of interest, the goal is to distinguish with high probability if $G$ satisfies the property $\mathcal{P}$, or whether it is “far” from all graphs that satisfy $\mathcal{P}$, with a suitable notion of farness. In [GGR98], the “dense” model was considered, where a graph is accessed through adjacency queries: for a pair of vertices $(u,v)$, the query reveals whether $(u,v)$ is an edge in the graph. In this model, a graph $G$ is $\varepsilon$-far from satisfying $\mathcal{P}$ if one needs to add or remove at least $\varepsilon N^{2}$ edges of $G$ to obtain a graph that satisfies $\mathcal{P}$.
In a later work, Goldreich and Ron [goldreich2002property] introduced the “bounded-degree” model for testing sparse graphs, focusing on the properties of bipartiteness and expansion. In this model, a $d$-bounded degree graph $G$ with $N$ vertices is accessed by performing neighbor queries: for a vertex $v$ and an integer $i\in[d]$, the query $(v,i)$ returns either the $i$-th neighbor of $v$, or some special symbol if $v$ has less than $i$ neighbors. The graph $G$ is said to be $\varepsilon$-far from some property $\mathcal{P}$, if one needs to add or delete at least $\varepsilon dN$ edges of $G$ to obtain a graph that satisfies $\mathcal{P}$. Over the last two decades, there has been a significant number of works in this model, and we refer the interested reader to the books by Goldreich [goldreich2017introduction] and Bhattacharyya and Yoshida [bhattacharyya2022property] and several surveys [Fischer01, DBLP:journals/fttcs/Ron09, Czumaj2010, DBLP:journals/siamdm/RubinfeldS11].
@@ 1.3 Related works on collision finding
The problem of collision finding is a ubiquitous problem in the field of algorithm theory with wide applications in cryptography. Here, given a sequence $s$ of $N$ integers, the goal is to find a duplicate in $s$. If one has the guarantee that $\Theta(N)$ elements of the sequence are duplicated, which is the case for instance when the sequence consists of uniformly random integers from $[N]$, it is well-known that classically $\Theta(\sqrt{N})$ queries are necessary and sufficient due to the birthday paradox. In the quantum model, this can be solved with query complexity $\Theta(N^{1/3})$ by the algorithm of Brassard, Høyer and Tapp [BHT]. The matching lower bound was first stated for a specific set of hard instances known as 22-to-11 (i.e., each integer appears exactly twice or not at all) by Aaronson and Shi [AaronsonShi]. For some constant integer $k\geq 3$, those results can be further extended to finding $k$-collisions in a random input with suitable alphabet size, so that it contains $\Theta(N)$ $k$-duplicates with high probability. The classical query complexity for this problem is $\Theta(N^{1-1/k})$ [hellweg2013property, peng2023optimal], and quantumly it is $\Theta\left(N^{\frac{1}{2}\left(1-\frac{1}{2^{k}-1}\right)}\right)$ [zhandry_collision]. The situation is more complex for non-random inputs.
Remarkably, the complexity of testing $k$-collision-freeness (i.e., the absence of $k$-collisions) is harder to settle on the lower bound side than the finding version. In this work, we are going to focus on the hardness of distinguishing inputs that have linearly many collisions from those that do not have any. For $k=2$, the two problems have the same complexity, since intuitively the only way to distinguish is to find a collision. This can be formalized easily in the classical case. Quantumly, this is more challenging, but the lower bound in [AaronsonShi] proved the hardness of distinguishing between 22-to-11 instances and ones with no duplicate.
@@ 1.4 Our results
Theorem 1.1 (Restated in Theorem˜3.3).
@@
Theorem 1.2 (Corollary of Theorem˜1.3).
@@
For proving graph property testing lower bounds, both the classical works of [hellweg2013property] and [peng2023optimal] prove collision testing lower bounds using the proportional moments technique of [raskhodnikova2009strong]. At the heart of this technique is a construction of two positive integer random variables, $X_{1}$ and $X_{2}$, with different expectations but with the following conditions on the first k1k-1 moments: $\mathop{\mathbb{E}}[X_{1}]/\mathop{\mathbb{E}}[X_{2}]=\mathop{\mathbb{E}}[X^{2}_{1}]/\mathop{\mathbb{E}}[X^{2}_{2}]=\ldots=\mathop{\mathbb{E}}[X^{k-1}_{1}]/\mathop{\mathbb{E}}[X^{k-1}_{2}]$. Such a construction leads to a randomized query complexity lower bound of $\Omega(N^{1-\frac{1}{k}})$ for various property testing problems such as $k$-collision-freeness [peng2023optimal]. Having a quantum version of this technique has been identified as an important open problem [ambainis2016efficient], since this could be used to pave the way to stronger quantum lower bounds in related settings. We modestly made progress to this quest for the special case of testing $k$-collision-freeness.
@@
Theorem 1.3 (Restated in Theorem˜4.1).
@@
Theorem 1.4 (Restated in Theorem˜5.1).
@@ 1.5.1 Subgraph-finding algorithm
We start by describing how to prove the upper bound result of Theorem˜1.1 for testing $k$-source-subgraph-freeness. We view the problem as a generalization of the problem of finding $k$-collisions and adapt an existing quantum algorithm for the latter problem. In [zhandry_collision], an algorithm is given for finding $k$-collisions in length-$N$ sequences of integers that contain $\Omega(N)$ $k$-collisions (e.g. $k$-to-11, or random sequence with appropriate parameters). Their algorithm generalizes the well-known collision finding algorithm of [BHT]. On a high level, the [zhandry_collision] algorithm first finds several $2$-collisions using Grover search like in [BHT], extends some of them to $3$-collisions in a similar way, and so on until a $k$-collision is found.
@@
When we make the second step of turning to testing of subgraph-freeness, we need to prove a variant of this statement: if an $N$-vertex graph $G$ is far from $H$-freeness (for some constant-sized subgraph $H$) then it contains $\Omega(N)$ many “source-disjoint” $H$-subgraphs. This means that there are $\Omega(N)$ many such $H$-subgraphs in $G$ that the set of vertices in the source components of each $H$-subgraph are disjoint. We prove this fact in Proposition˜3.2 and this allows us to further generalize the approach of [zhandry_collision]: first find several partial solutions where only a few source components of an $H$-subgraph are explored, and gradually extend these (using Grover search coupled with constant-depth BFS) until a complete $H$-subgraph is found.
@@ 1.5.2 Collision-freeness lower bound
Now we will discuss our approach to proving the lower bounds of collision-freeness (Theorem˜1.2) and $k$-source-subgraph-freeness (Theorem˜1.3). We first give a simple reduction from $k$-collision-freeness to $k$-star-freeness, which is a special case of $k$-source-subgraph-freeness. This way, it is enough to prove a lower bound on testing $k$-collision-freeness, and it implies the same result on testing $k$-source-subgraph-freeness. Since our lower bound approach crucially depends on the (dual) polynomial method, let us start by briefly discussing it.
@@
The polynomial method can be written in the form of a linear program, of which one can take the dual. By weak LP-duality, when using this dual characterization for proving a lower bound on function $f$, one needs to provide a “witness” of the approximating polynomial’s high degree, say $\Delta$. This witness is called the dual polynomial $\psi$ and, in the easiest case of total Boolean [^2] We use $\{-1,1\}$ where 1-1 corresponds to the “true value”. functions $f:\{-1,1\}^{n}\to\{-1,1\}$, it needs to have three properties:
@@
    Pure high degree $\Delta$: $\sum_{x}p(x)\psi(x)=0$, for every polynomial $p$ with degree $\Delta$,
@@
[…]
@@
Moreover, the symmetries of $f$ allow the extension of the initial function $f$ from the very restricted set of binary inputs correspond to valid strings, to the more general set of binary inputs of Hamming weight $N$ [ambainis2016efficient]. With further technicalities one can also extend $f$ to all binary inputs of Hamming weight at most $N$ [Bun_Thaler_20]. This is fundamental because instead of being forced to zero out the dual polynomial outside the domain of $f$, we only need to do so on inputs of Hamming weight higher than $N$. Using the symmetry of $f$, it can be shown that a lower bound on this modified, Boolean version implies a lower bound on the original $k$-collision problem.
@@
    Pure high degree $\Delta$: $\sum_{x}p(x)\psi(x)=0$, for every polynomial $p$ with degree $\Delta$;
@@
Coming back now to the definition of our Boolean function $f$, one can rewrite it as a composition of simpler functions: $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$, where by composition we mean $(g\circ h)(x)=g(h(x_{1}),\dots,h(x_{n}))$ (where $x=(x_{1},\dots,x_{n})$ and each $x_{i}$ is a binary vector of appropriate dimension) and the domain is restricted to bit strings of Hamming weight at most $N$. Note that here, $x_{j}=(x_{1,j},x_{2,j},\ldots,x_{N,j})$. Above, $\mathrm{THR}_{N}^{k}$ is the threshold function: it is 1-1 if the input bitstring contains at least $k$ many 1-1 (true) values, and is $1$ otherwise; and $\mathrm{GapOR}_{R}^{\gamma}$ is the gap version of $\mathrm{OR}$, which is $1$ if the input is all-1, 1-1 if the input contains at least $\gamma R$ many 1-1 values, and is undefined otherwise.
In order to give a dual polynomial for this composed function, we start from a dual polynomial for each part of the composition ($\phi$ and $\psi$), which were already given in [polynomial_strikes_back] (in different contexts). Then we use a known way [ShiZ09, Lee09, Sher13] of composing dual polynomials called the dual block composition, which provides a nearly good dual polynomial $\phi\star\psi$ for $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$. Indeed, by construction, the normalization (ii) and the pure high degree (iii) are guaranteed. However, the issue is that it is not 0 on bitstrings of large Hamming weight thus (iv) is not satisfied, and and the high correlation (i’) still has to be proved.
To fix (iv), we use another result of [polynomial_strikes_back] which provides another dual polynomial $\zeta$, that is close to $\phi\star\psi$ and that is 0 on inputs having Hamming weight larger than $N$. Also, it only changes the pure high degree by a polylogarithmic factor. Now the only remaining task is to prove a large enough correlation (i’) of $\phi\star\psi$ and $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$ so that $\zeta$ still has high enough correlation. This high correlation proof (Lemma˜4.18) is the most technical part of our paper.
High correlation: proof of Lemma˜4.18
@@
In the proof of this lemma we use Proposition˜4.17 that is a more general statement of some techniques used in several proofs of [polynomial_strikes_back]. But then we need to diverge from their proof because it crucially relies on a certain one-sided error property (in the sense of [polynomial_strikes_back, Lemma 6.11]) of the inner function of the composition which is OR function in their case. Our inner function is the threshold function which does not satisfy this property, so we have to use some other properties of the dual polynomials in our proof. This different proof technique could be a step towards obtaining a more general lower bound technique.
@@
Let us now discuss our approach to proving the linear lower bound on quantum query complexity of testing $3$-colorability (Theorem˜1.4). Before proceeding to present our approach, let us briefly discuss the classical lower bound of testing $3$-colorability. To prove the classical lower bound of $3$-colorability, the authors in [bogdanov2002lower] first studied another problem called E(3,c)(3,c)LIN-2, a problem related to deciding the satisfiability of a system of linear equations. More formally, E(3,c)(3,c)LIN-2 considers a system of linear equations modulo $2$, where each equation has $3$ variables and every variable appears in at most $c$ equations. Given such a system of linear equations, the goal is to distinguish if it is satisfiable, or at least some suitable fraction of the equations need to be modified to satisfy it. [bogdanov2002lower] proved that $\Omega(N)$ classical queries to the system of linear equations are necessary for testing E(3,c)(3,c)LIN-2.
After this, they designed a reduction from E(3,c)(3,c)LIN-2 to $3$-colorability such that satisfying instances of E(3,c)(3,c)LIN-2 are reduced to $3$-colorable graphs, and far from satisfiable instances of E(3,c)(3,c)LIN-2 are mapped to far from $3$-colorable graphs. Combining these two arguments, the authors in [bogdanov2002lower] proved that $\Omega(N)$ classical queries are necessary for testing $3$-colorability for bounded degree graphs.
The authors in [bogdanov2002lower] used Yao’s minimax method to prove the linear lower bound in testing E(3,c)(3,c)LIN-2. In particular, they designed two distributions $D_{\mathrm{yes}}$ and $D_{\mathrm{no}}$ such that the systems of linear equations in $D_{\mathrm{yes}}$ are satisfiable, whereas the systems of linear equations in $D_{\mathrm{no}}$ are far from being satisfiable. A crucial ingredient of their lower bound proof is a construction of a system of linear equations (represented as a matrix) that are far from being satisfiable, but any $\delta N$ rows of the matrix are linearly independent. Hence, any subset of $\delta N$ entries of the matrix will look uniformly random, and therefore hard to distinguish from a satisfiable instance.
It is a known fact that distinguishing between a uniformly random string and a $\ell$-wise independent string, for an appropriate integer $\ell$, is hard for quantum algorithms (see e.g. [apers2022quantum]). Using this result, we can construct suitable hard instances for E(3,c)(3,c)LIN-2 such that testing E(3,c)(3,c)LIN-2 remains maximally hard (requires $\Omega(N)$ queries) for any quantum algorithm. Combining this hardness result with the reduction from E(3,c)(3,c)LIN-2 to $3$-colorability, we finally prove that $\Omega(N)$ quantum queries are necessary for testing $3$-colorability. We formally prove this in Section˜5.